OpenStudy (danysanchy97):

A bus on a regular schedule takes 3 1/4 hours to reach its destination. The express bus takes 2 1/2 hours to make the same trip. How much travel time can be saved by taking the express? A. 3/4 hour B. 5 3/4 hours C. 2 hours D. 1 1/4 hours

OpenStudy (boldjon):

3-1/4 - 2-1/2 = ? convert to lowest common denominator: 3-1/4 - 2-2/4 = ? get rid of mixed fractions and solve (3*4)+1 = 13 (2*4)+2 = 10 13/4 - 10/4 = 3/4 A.) 3/4 hours

OpenStudy (danysanchy97):

thank you. can you help me on some more?